find the product of -i,2i and (-1/8i)power3 pleace get answer bros

Grade:11

2 Answers

Vijay Mukati
8 years ago
Dear Student,

The product of -i * 2i * (-1/8i)3 = 2i2/8i3 = 1/4i = -i/4.
Just Remember i2= -1

Thanks.
siva prasad reddy
8 years ago
= -12 = 1/4i = -i/4.Just Remember i3/8i2 = 2i3The product of -i * 2i * (-1/8i)
thanks.....................................................................................................